# On pro-cdh descent on derived schemes

### v1

@@ 6. Generalised Weibel vanishing
The following theorem generalizes [KST18b, Thm. B], which treats the case of Noetherian classical schemes.
@@
Let $X$ be a qcqs derived scheme. Assume that the underlying topological space of the spectrum of every local ring of every irreducible component of $\prescript{\mathrm{cl}}{}{X}$ is Noetherian. Then the following hold.
@@
Remarks 6.2.
  1. (1)
    We point out that Lemma 6.6 on the annihilation of negative $K$-theory classes is the only ingredient in our proof of Theorem 6.1 where we need a Noetherian assumption on the underlying topological space.
  2. (2)
    Any qcqs scheme of Krull dimension $\leq 1$ satisfies the assumptions of the theorem.
  3. (3)
    The assumptions in the theorem are stable under maps essentially of finite type: Assume that $X$ satisfies the hypothesis of Theorem 6.1. If $f\colon Y\to X$ is a morphism of qcqs derived schemes such that the morphism $\prescript{\mathrm{cl}}{}{f}$ of underlying classical schemes is essentially of finite type, then also $Y$ satisfies the hypothesis of the theorem.
    Indeed, we may assume that $Y$ and $X$ are classical. Let $Y_{0}$ be the spectrum of a local ring of an irreducible component of $Y$. Then the restriction of $f$ to $Y_{0}$ factors through the spectrum $X_{0}$ of a local ring of an irreducible component of $X$ and the induced morphism $Y_{0}\to X_{0}$ is essentially of finite type. Say $X_{0}=\operatorname{Spec}(A)$ and $Y_{0}=\operatorname{Spec}(S^{-1}B)$ where $B$ is a finitely generated $A$-algebra and $S\subseteq B$ is a multiplicative subset. As the topological space $|X_{0}|$ is Noetherian by assumption, so is $|\operatorname{Spec}(B)|$ by [OP68, Cor. 2.6]. As $\operatorname{Spec}(S^{-1}B)$ is homeomorphic to a subspace of $\operatorname{Spec}(B)$, it is Noetherian, too.
For proving Theorem 6.1, we follow the strategy of the proof of [KST18b, Thm. B] though we need some additional inputs. The following lemma replaces [KS17, Lemma 4] and is essentially due to Scheiderer [Sch92].
@@
For a scheme $X$ and a point $x\in X$ with closure $\overline{\{x\}}$ in $X$,
In particular,
Proof.
By definition it suffices to prove the lemma in case $X=\operatorname{Spec}(A)$ with $A$ integral. In this case, the first assertion is [Jaf60, Ch. IV, Prop. 2], the second assertion follows from the fact that $\dim(Y)\leq\operatorname{vdim}(Y)$ for every scheme $Y$. ∎
Lemma 6.5.
Let $X$ be a qcqs derived scheme of finite valuative dimension. Let $F$ be a sheaf of spectra on $X$ for the Zariski topology. Assume that for every $x\in X$, the homotopy groups of the stalks $\pi_{-i}(F_{x})$ vanish for $i>\operatorname{vdim}(\mathcal{O}_{X,x})$. Then $\pi_{-i}(F(X))=0$ for all $i>\operatorname{vdim}(X)$.
@@
In the following lemma, given a morphism $f\colon Y\to X$, we abusively denote the induced morphism $\mathbb{A}^{r}_{Y}\to\mathbb{A}^{r}_{X}$ by the same letter.
Lemma 6.6.
Let $X=\operatorname{Spec}(A)$ be a classical integral affine scheme whose underlying topological space is Noetherian. Let $i>0$, $r\geq 0$, and let $\gamma$ be an element in $K_{-i}(\mathbb{A}^{r}_{X})$. Then there exists an irreducible derived scheme $Y$ and a proper lafp morphism $f\colon Y\to X$ which is an isomorphism over a non-empty open subset $U\subseteq X$ with the property that $f^{*}(\gamma)=0$ in $K_{-i}(\mathbb{A}^{r}_{Y})$.
Proof.
As $K$-theory commutes with filtered colimits of rings, there exists a subring $A_{0}\subseteq A$, finitely generated as a $\mathbb{Z}$-algebra, and an element $\gamma_{0}\in K_{-i}(\mathbb{A}^{r}_{X_{0}})$ with $a^{*}(\gamma_{0})=\gamma$. Here $X_{0}=\operatorname{Spec}(A_{0})$ and $a\colon X\to X_{0}$ is the induced morphism.
By [KS17, Prop. 5] there exists a proper birational morphism of schemes $f_{0}\colon Y_{0}\to X_{0}$ such that $f_{0}^{*}(\gamma_{0})=0$ in $K_{-i}(\mathbb{A}^{r}_{Y_{0}})$. As $X_{0}$ is Noetherian, $f_{0}$ is lafp (see 2.4). Let $f_{1}\colon Y_{1}\to X$ be the derived pullback of $f_{0}$ along $a\colon X\to X_{0}$. Then $f_{1}$ is proper, lafp, and an isomorphism over a non-empty open subscheme $U\subseteq X$. From the commutative diagram
[…]
it is clear that $f_{1}^{*}(\gamma)=0$ in $K_{-i}(\mathbb{A}^{r}_{Y_{1}})$.
As $\prescript{\mathrm{cl}}{}{f}_{1}$ is of finite type, the topological space $|Y_{1}|$ is Noetherian [OP68, Cor. 2.6]. In particular, the open complement of the closure of $|U|$ in $|Y_{1}|$ is quasi-compact, and hence there exists a classically finitely presented closed subscheme $Y_{2}\hookrightarrow\prescript{\mathrm{cl}}{}{Y}_{1}$ whose underlying topological space $|Y_{2}|$ is the closure of $|U|$ in $|Y_{1}|$ (see footnote 4 in the proof of Theorem 3.2 or [KS24, Lemma 4.14(1)] for an alternative argument). Note that $Y_{2}\hookrightarrow Y_{1}$ is an isomorphism over $U$. By Lemma 2.2 there exists a finitely presented closed subscheme $Y\hookrightarrow Y_{1}$ with $\prescript{\mathrm{cl}}{}{Y}=Y_{2}$. By construction, $Y$ is irreducible, the composite $f\colon Y\hookrightarrow Y_{1}\to X$ is proper, lafp, and an isomorphism over $U$, and $f^{*}(\gamma)=0$, as desired. ∎
In the proof, we use the following well known facts about non-positive $K$-theory on affine (derived) schemes.
@@
(1) For connective $K$-theory, this is due to Waldhausen [Wal78, Prop. 1.1]. The general case follows from this together with [BGT13, Thm. 9.53] (or [KST18b, Thm. 2.16]). Alternatively, see [LT19, Lemma 2.4] for a slightly more general statement.
@@
We prove the theorem by induction on $d=\operatorname{vdim}(X)$. By Lemma 6.5, applied to $K$ and $\Sigma N^{(r)}K$ respectively, we may assume that $X$ is affine and local. In this case, $K_{-i}(\mathbb{A}^{r}_{X})\cong K_{-i}(\mathbb{A}^{r}_{\prescript{\mathrm{cl}}{}{X}^{\mathrm{red}}})$ for all $i\geq 0$ and all $r\geq 0$ by Lemma 6.7 (1), (2). So we can assume that $X$ is a classical reduced affine local scheme.
@@
Now assume that $d$ is positive and the assertion is proven for all derived schemes satisfying the assumptions of Theorem 6.1 that are of valuative dimension $<d$. We first prove the result for $X$ under the additional assumption that $X$ is irreducible, and hence has Noetherian topological space.
Let $\gamma$ be an element in $K_{-i}(X)$ with $i>\operatorname{vdim}(X)$ or in $N^{(r)}K_{-i}(X)$ with $i\geq\operatorname{vdim}(X)$. We have to show that $\gamma=0$. Let $f\colon Y\to X$ be a proper, lafp morphism as provided by Lemma 6.6. In particular $f^{*}(\gamma)=0$. Let $Z$ be the closed complement of $U$ in $X$. By Theorem 4.5 we obtain an exact sequence of pro-abelian groups
$K_{-i+1}(Y_{Z}^{\wedge})\to K_{-i}(X)\to K_{-i}(Y)\oplus K_{-i}(X_{Z}^{\wedge})$
and similarly for $N^{(r)}K$ in place of $K$. Recall from (1) in 2.5 that $K_{-i}(X_{Z}^{\wedge})$ is the pro-abelian group $\{K_{-i}(Z^{\prime})\}_{Z^{\prime}}$ where $Z^{\prime}$ runs through all closed (derived) subschemes $Z^{\prime}\hookrightarrow X$ with $|Z^{\prime}|=Z$ and similarly $K_{-i+1}(Y_{Z}^{\wedge})=\{K_{-i+1}(Y\times_{X}Z^{\prime})\}_{Z^{\prime}}$.
As $Z$ does not contain the generic point of $X$, we have $\operatorname{vdim}(Z^{\prime})<\operatorname{vdim}(X)$ for every $Z^{\prime}$ as above by [EHIK20, Prop. 2.3.2(4)]. Hence $K_{-i}(X_{Z}^{\wedge})$ vanishes by induction. By [EHIK20, Prop. 2.3.2(6)] we have $\operatorname{vdim}(Y)=\operatorname{vdim}(X)$. As $f^{-1}(Z)$ does not contain the generic point of $Y$, the same argument as before implies that $\operatorname{vdim}(Y\times_{X}Z^{\prime})<\operatorname{vdim}(Y)$ for all $Z^{\prime}$ as above. Hence also $K_{-i+1}(Y_{Z}^{\wedge})$ vanishes by induction. It follows that $f^{*}\colon K_{-i}(X)\to K_{-i}(Y)$ is injective. Similarly, $f^{*}\colon N^{(r)}K_{-i}(X)\to N^{(r)}K_{-i}(Y)$ is injective. As $f^{*}(\gamma)=0$, we must have $\gamma=0$, as desired.
We now consider the general case. So $X$ is now a classical, reduced, affine, local scheme of valuative dimension $d>0$ and every irreducible component of $X$ has Noetherian topological space. For ease of notation, we only consider assertion (1) of the theorem. The proof of (2) is completely parallel. To reduce to the integral case treated above, we apply an argument from [EHIK20, Thm. 2.4.15] as follows: Take an integer $i>d=\operatorname{vdim}(X)$ and an element $\gamma\in K_{-i}(X)$. We wish to show that $\gamma=0$. Let
@@
Proof.
We note that homotopy $K$-theory is a cdh sheaf. This was first proven by Cisinski [Cis13] for Noetherian schemes of finite dimension, which implies the general statement by absolute Noetherian approximation; alternatively it follows from the fact that $KH$ is truncating, see [LT19, Cor. A.5]. In particular, we get the maps

### v2

@@ 6.3. Weibel vanishing
The following theorem generalizes [KST18, Thm. B], which treats the case of Noetherian classical schemes.
@@
Let $X$ be a qcqs spectral scheme. Then the following hold.
@@
In the spectral setting, the affine space in (2) is either the flat one or the smooth one, the assertion holds for both. In fact, one reduces immediately to underlying classical schemes (see the beginning of the proof in Subsection 6.5), and in both cases this is just the classical affine space.
The proof of Theorem 6.5 is along the lines of Kerz’s proof in the Noetherian case in [Ker18] which is a bit more direct than the one in [KST18, Thm. B]. It uses the theory of reductions presented in Subsection 6.2 in order to prove better descent results for blowups. The following lemma replaces [KS17, Lemma 4] and is essentially due to Scheiderer [Sch92].
@@
Let $X$ be a qcqs spectral scheme of finite valuative dimension. Let $F$ be a sheaf of spectra on $X$ for the Zariski topology. Assume that for every $x\in X$, the homotopy groups of the stalks $\pi_{-i}(F_{x})$ vanish for $i>\operatorname{vdim}(\mathcal{O}_{X,x})$. Then $\pi_{-i}(F(X))=0$ for all $i>\operatorname{vdim}(X)$.
@@
6.4. $K$-theoretic preparations
In the proof of Theorem 6.5, we use the following well known facts about non-positive $K$-theory on affine (derived) schemes.
@@ 6.4. $K$-theoretic preparations
(1) For connective $K$-theory, this is due to Waldhausen [Wal78, Prop. 1.1]. The general case follows from this together with [BGT13, Thm. 9.53] (or [KST18, Thm. 2.16]). Alternatively, see [LT19, Lemma 2.4] for a slightly more general statement.
@@
The previous lemma has as a simple consequence the following lemma which will allow us to get some control on the negative $K$-theory of certain blowups in Proposition 6.12.
Lemma 6.9.
Let $X$ be a separated derived scheme admitting a covering by $\ell$ affine open subschemes. Then the canonical morphism
$K(X)\to K(\prescript{\mathrm{cl}}{}{X}^{\prime}),$
where $\prescript{\mathrm{cl}}{}{X}^{\prime}$ denotes $\prescript{\mathrm{cl}}{}{X}$ with any closed subscheme structure between $\prescript{\mathrm{cl}}{}{X}^{\mathrm{red}}$ and $\prescript{\mathrm{cl}}{}{X}$, is $(-\ell+2)$-connective.
Proof.
Induction on $\ell$. If $\ell=1$, this is Lemma 6.8 (1, 2). For $\ell>1$ write $X=V\cup U$ with $U$ affine and $V$ admitting an open covering by $\ell-1$ affines. As $X$ is separated, also $V\cap U$ has an open covering by $\ell-1$ affines. The claim follows from the inductive hypothesis by considering the map from the cartesian square
${K(X)}$${K(V)}$${K(U)}$${K(V\cap U)}$
to the similar cartesian square with the underlying classical schemes with the subscheme structure induced from $X^{\prime}$. ∎
In the following, if $X$ is a qcqs derived scheme and $Z\hookrightarrow X$ is a derived subscheme, we denote by $K(X,Z)$ the relative $K$-theory $\operatorname{fib}(K(X)\to K(Z))$.
Lemma 6.10.
Let $X$, $Y$ be quasi-compact, separated classical schemes, $Z\hookrightarrow X$ a classically finitely presented closed subscheme. Let $f\colon Y\to X$ be a finite morphism which is an isomorphism outside $Z$. Assume that $X$ has a covering by $\ell$ affine open subschemes. Then the canonical map
$K(X,Z)\to K(Y,f^{-1}(Z))$
is $(-\ell+1)$-connective.
Proof.
By induction on $\ell$, as in the proof of Lemma 6.9, it suffices to treat the case $\ell=1$, i.e. $X$ and hence $Y$ affine. As $K$-theory commutes with filtered colimits, by Lemma 6.11 below we may assume that $f\colon Y\to X$ is finitely presented. By Remark 4.3 the square of pro-spectra
${K(X)}$${K(X_{Z}^{\wedge})}$${K(Y)}$${K(Y_{Z}^{\wedge})}$
is then weakly cartesian. Hence the square of spectra obtained by applying $\lim\colon\operatorname{Pro}(\operatorname{Sp})\to\operatorname{Sp}$,
${K(X)}$${\lim K(X_{Z}^{\wedge})}$${K(Y)}$${\lim K(Y_{Z}^{\wedge}),}$
is cartesian. By Lemma 6.8(1, 2), each transition map in the pro-spectrum $K(X_{Z}^{\wedge})$ is 1-connective. In particular, using the Milnor sequence we get isomorphisms $\pi_{0}(\lim K(X_{Z}^{\wedge}))\cong\lim K_{0}(X_{Z}^{\wedge})\cong K_{0}(Z)$ and surjections $\pi_{1}(\lim K(X_{Z}^{\wedge}))\to\lim K_{1}(X_{Z}^{\wedge})\to K_{1}(Z)$. In other words, $\lim K(X_{Z}^{\wedge})\to K(Z)$ is 1-connective. Consider the following diagram in which the left vertical map is an equivalence by the previous cartesian square.
[…]
The fibre of the top horizontal map identifies with $\Omega\operatorname{fib}(\lim K(X_{Z}^{\wedge})\to K(Z))$ and is hence 0-connective. By the same arguments, the lower horizontal map is 0-connective. It follows that also the right vertical map is 0-connective. ∎
Lemma 6.11.
Let $\varphi\colon A\to B$ be a finite morphism of rings such that $\operatorname{Spec}(\varphi)\colon\operatorname{Spec}(B)\to\operatorname{Spec}(A)$ is an isomorphism outside $V(I)$ where $I=(f_{1},\dots,f_{r})\subseteq A$. Then we can write $\varphi$ as a filtered colimit of finitely presented, finite maps $\varphi_{\lambda}\colon A\to B_{\lambda}$ such that each $\operatorname{Spec}(\varphi_{\lambda})$ is an isomorphism outside $V(I)$.
Proof.
Choose $A$-module generators $b_{1},\dots,b_{s}$ of $B$. We get a surjection $A[T_{1},\dots,T_{s}]\to B$, $T_{i}\mapsto b_{i}$. Let $J$ denote its kernel. As $b_{i}$ is integral over $A$, there exists a monic polynomial $p_{i}\in A[X]$ with $p_{i}(b_{i})=0$. Clearly, $p_{i}(T_{i})\in J$.
Let $f$ be one of the generators $f_{j}$ of $I$. As $\operatorname{Spec}(\varphi)$ is an isomorphism outside $V(I)$, we have $A[f^{-1}]\xrightarrow{\simeq}B[f^{-1}]$. In particular, $B[f^{-1}]$ is finitely presented as an $A[f^{-1}]$-algebra, and it follows that $J[f^{-1}]\subseteq A[f^{-1}][T_{1},\dots,T_{s}]$ is a finitely generated ideal. We may pick finitely many generators $q_{jk}$ which already live in $J$.
Now write $J$ as increasing union of finitely generated ideals $J_{\lambda}$ where each $J_{\lambda}$ contains all the $p_{i}(T_{i})$ and all $q_{jk}$’s. By construction, $B_{\lambda}=A[T_{1},\dots,T_{s}]/J_{\lambda}$ is then finitely presented over $A$, finite (as in $B_{\lambda}$ each $T_{i}$ is integral over $A$), and $\operatorname{Spec}(B_{\lambda})\to\operatorname{Spec}(A)$ is an isomorphism outside $V(I)$ (as $J_{\lambda}[f_{j}^{-1}]=J[f_{j}^{-1}]$ for all $j$). Clearly, $\operatorname*{colim}_{\lambda}B_{\lambda}=B$. ∎
Proposition 6.12.
Let $R$ be a local ring, $I\subseteq R$ be a finitely generated, proper ideal of analytic spread $\ell=\ell(I)$. Let $Z=V(I)$, $f\colon\operatorname{Bl}_{Z}(X)\to X$ the blowup, and $E=f^{-1}(Z)$ the classical exceptional divisor. Then the pullback map
$K(X,Z)\to K(\operatorname{Bl}_{Z}(X),E)$
is $(-\ell+1)$-connective. More generally, for any $r\geq 0$, the pullback map
$K(\mathbb{A}^{r}\times X,\mathbb{A}^{r}\times Z)\to K(\mathbb{A}^{r}\times\operatorname{Bl}_{Z}(X),\mathbb{A}^{r}\times E)$
is $(-\ell+1)$-connective.
Proof.
By Proposition 6.4, there exists a reduction $J$ of some power $I^{n}$ of $I$ generated by $\ell$ elements, say $J=(f_{1},\dots,f_{\ell})$. As $\operatorname{Bl}_{I^{n}}(R)=\operatorname{Bl}_{I}(R)$ canonically, Lemma 6.2 yields a finite $X$-morphism $\operatorname{Bl}_{Z}(X)=\operatorname{Bl}_{I}(R)\to\operatorname{Bl}_{J}(R)$. Let $\widetilde{X}$ be the derived blowup of $X$ in $f_{1},\dots,f_{\ell}$. Then we have a closed immersion $\operatorname{Bl}_{J}(R)\hookrightarrow\widetilde{X}$ and in total a finite morphism $\operatorname{Bl}_{I}(X)\to\widetilde{X}$. Moreover, $\widetilde{X}$ has a covering by $\ell$ affine open subschemes. Write $D=\prescript{\mathrm{cl}}{}{(}\widetilde{X}\times_{X}Z)$. We now argue similarly as in the proof of Lemma 6.10. Consider the following commutative diagram.
[…]
The left vertical map is an equivalence by Proposition 4.1 (note that $|Z|=|V(J)|$). The top horizontal map is 0-connective, the lower horizontal map is $(-\ell+1)$-connective by Lemma 6.9. Hence the right vertical map is $(-\ell+1)$-connective. Using Lemma 6.9 again, we see that $K(\widetilde{X},D)\to K(\prescript{\mathrm{cl}}{}{\widetilde{X}},D)$ is $(-\ell+2)$-connective. In total, $K(X,Z)\to K(\prescript{\mathrm{cl}}{}{\widetilde{X}},D)$ is $(-\ell+1)$-connective. [^5] In fact, using the above connectivity estimates, this follows more easily from the derived generalization of Thomason’s theorem [Tho93] in [KST18, Thm. 3.7] which holds without any Noetherianity assumptions (with the same proof as in loc. cit.) By Lemma 6.10, the map $K(\prescript{\mathrm{cl}}{}{\widetilde{X}},D)\to K(\operatorname{Bl}_{Z}(X),E)$ is $(-\ell+1)$-connective. Hence the composite is $(-\ell+1)$-connective, as asserted.
The more general claim can be proved by exactly the same argument, using that $\mathbb{A}^{r}$ is affine and flat over $\mathbb{Z}$ and derived blowups commute with arbitrary base change. ∎
6.5. Proof of Weibel vanishing
@@
We may assume that $d=\operatorname{vdim}(X)$ is finite and prove the theorem by induction on $d$. By Lemma 6.7, applied to $K$ and $\Sigma N^{(r)}K$ respectively, we may assume that $X$ is affine and local. In this case, $K_{-i}(\mathbb{A}^{r}_{X})\cong K_{-i}(\mathbb{A}^{r}_{\prescript{\mathrm{cl}}{}{X}^{\mathrm{red}}})$ for all $i\geq 0$ and all $r\geq 0$ by Lemma 6.8 (1), (2). So we can assume that $X$ is a classical reduced affine local scheme.
@@
Now assume that $d$ is positive and the assertion is proven for all qcqs derived schemes that are of valuative dimension $<d$.
Claim.
The assertion of Theorem 6.5 holds for $X$ as above under the additional assumption that $X$ is irreducible.
Proof of the claim..
Write $X=\operatorname{Spec}(R)$. So $R$ is a local integral domain. Let $\gamma$ be an element in $K_{-i}(R)$ with $i>d$ or in $N^{(r)}K_{-i}(R)\subseteq K_{-i}(\mathbb{A}^{r}_{R})$ with $i\geq d$ in case $r>0$. We have to show that $\gamma=0$.
As $K$-theory commutes with filtered colimits, there exists a subring $R_{0}$ of $R$, finitely generated as a $\mathbb{Z}$-algebra, and $\gamma_{0}\in K_{-i}(\mathbb{A}^{r}_{R_{0}})$ pulling back to $\gamma$. By [KS17, Prop. 5], there exists a (finitely generated) ideal $I_{0}\subseteq R_{0}$ such that $\gamma_{0}$ is annihilated in $K_{-i}(\mathbb{A}^{r}\times\operatorname{Bl}_{I_{0}}(R_{0}))$. Let $I=RI_{0}\subseteq R$, and $Z=V(I)$. By functoriality of the blowup, $\gamma$ is annihilated in $K_{-i}(\mathbb{A}^{r}\times\operatorname{Bl}_{Z}(X))$. Let $E\subseteq\operatorname{Bl}_{Z}(X)$ denote the exceptional divisor. By construction $Z\subseteq X$ is a finitely presented, proper subscheme, hence $\operatorname{Bl}_{Z}(X)\to X$ is a modification and thus $\operatorname{vdim}\operatorname{Bl}_{Z}(X)=\operatorname{vdim}X=d$ [EHIK20, Prop. 2.3.2(6)]. Furthermore $\operatorname{vdim}(Z),\operatorname{vdim}(E)<d$ by loc. cit., Assertion 4.
We now treat the case of $\gamma\in K_{-i}(R)$ with $i>d$. Consider the commutative diagram of exact sequences
${K_{-i+1}(Z)}$${K_{-i}(X,Z)}$${K_{-i}(X)}$${K_{-i}(Z)}$${K_{-i+1}(E)}$${K_{-i}(\operatorname{Bl}_{Z}(X),E)}$${K_{-i}(\operatorname{Bl}_{Z}(X))}$${K_{-i}(E).}$
As $\operatorname{vdim}(Z),\operatorname{vdim}(E)<d$, the four outer groups vanish by induction. If $I=R$, then $\operatorname{Bl}_{Z}(X)=X$ and hence $\gamma=0$. So we may now assume that $I\subset R$ is a proper ideal. By Proposition 6.3 we then have $\ell(I)\leq\operatorname{vdim}(R)=d$. As $i>d$, Proposition 6.12 implies that the second vertical map is an isomorphism. Hence $K_{-i}(X)\to K_{-i}(\operatorname{Bl}_{Z}(X))$ is an isomorphism, too, and hence $\gamma=0$.
The argument in the case $\gamma\in N^{(r)}K_{-i}(R)$ with $i\geq d$ is exactly the same with the $K$-groups above replaced by the $N^{(r)}K$-groups. This finishes the proof of the Claim. ∎
We now consider the general case. So $X$ is now a classical, reduced, affine, local scheme of valuative dimension $d>0$. For ease of notation, we only consider assertion (1) of the theorem. The proof of (2) is completely parallel. To reduce to the integral case treated in the Claim above, we apply an argument from [EHIK20, Thm. 2.4.15] as follows: Take an integer $i>d=\operatorname{vdim}(X)$ and an element $\gamma\in K_{-i}(X)$. We wish to show that $\gamma=0$. Let
@@
Proof of Theorem 1.4.
We note that homotopy $K$-theory is a cdh sheaf. This was first proven by Cisinski [Cis13] for Noetherian schemes of finite dimension (and in characteristic 0 previously by Haesemeyer [Hae04]), which implies the general statement by absolute Noetherian approximation; alternatively it follows from the fact that $KH$ is truncating, see [LT19, Cor. A.5]. In particular, we get the maps